# Periodic Cosmic String Formation and Dynamics

### v1

@@ 1 Introduction and Summary
The formation of topological defects in theories with global or gauged $U(1)$ symmetries has been extensively studied in both high-energy and condensed-matter contexts Kibble:1976sj ; Kibble:1980mv ; Zurek:1996sj . The Kibble–Zurek mechanism Kibble:1976sj ; Zurek:1993ek ; Zurek:1985qw is a well-known example in thermal systems, in which topological defects form as the early Universe cools below the critical temperature for spontaneous symmetry breaking. Formation of cosmic strings can also occur non-thermally as a result of quantum fluctuations during inflation Linde:1990yj ; PhysRevD.44.340 ; Vishniac:1986sk . After inflation, nonlinearities may drive the formation of topological defects in a large background field Abrikosov:1956sx ; osti_4551407 ; Galaiko1966FormationOV ; East:2022ppo ; East:2022rsi ; Kofman:1997yn . In the gauged-$U(1)$ case (i.e., the Abelian Higgs model), in a system where the gauge field strength grows East:2022ppo ; East:2022rsi (either in the early Universe or via black-hole superradiance) and surpasses a critical threshold, a superheated phase transition occurs and a dense network of strings is produced. [^1] We will use the terms ‘vortices’ and ‘strings’ when discussing the dynamics presented in this paper. Although vortices are generally associated with $(2+1)$-dimensional systems and strings with $(3+1)$-dimensional systems, we use them interchangeably, depending on the context. For all practical purposes of this paper, there is no difference between these two terms. This string network absorbs energy from the background gauge fields, grows in size, and persists in the system. These dynamics resemble the vortex-forming phase transition in a superconductor, which is described in the framework of the Ginzburg–Landau model Ginzburg:1950sr ; Abrikosov:1956sx (see App. A for a short review of the relevant literature).
@@
The paper is organized as follows. In Sec. 2, we begin with a purely analytical perturbative analysis valid for small $U(1)$ charge density that demonstrates the existence of a parametric resonance instability Tkachev:1998dc for spatial field perturbations $\delta\Phi$ around a spatially homogeneous background field evolution $\Phi_{0}$ [Secs. 2.1 and 2.2]. We extend this perturbative analysis to larger $U(1)$ charge densities via a numerical study of the linearized system of equations governing the evolution of the perturbations $\delta\Phi$ Co:2020dya , demonstrating that this parametric resonance remains present in this case [Sec. 2.3]. We then analyze the fully non-linear evolution of the $\Phi$ field from the epoch of the initial parametric resonance through to the onset of the formation of the counter-rotating regions that emerge when the perturbations grow beyond the linear regime phase-coherently over a range of momentum modes [Sec. 2.4]; we do this both for small and, for the first time, large $U(1)$ charge densities. Finally, we summarize the main points of this section [Sec. 2.5].
@@ 2 Perturbation Growth to the Onset of Vortex Formation
In this section we show that a spatially homogeneous scalar field subject to a wine-bottle potential, initially either with or without field-space angular momentum, is not a stable field configuration: rather, it exhibits (classical) parametric instabilities that lead to growth of spatial inhomogeneities. Some of the results presented in this section were previously discussed in, e.g., Kofman:1994rk ; Co:2020dya .
For now, we ignore Hubble expansion and work in $(D+1)$-dimensional Minkowski spacetime (mostly minus convention; signature $1-D$). The Lagrangian (1) yields the equations of motion (EoM)
@@ 2.1 Background Field Evolution
Although the background-field trajectory $\Phi_{0}(t)$ is completely specified by a choice of the two conserved quantities $\mathcal{J}_{0}$ and $\mathcal{E}_{0}$, supplemented with an initial field point on the trajectory fixed by $\rho_{0}(0)\equiv\rho_{i}$ and $\theta_{0}(0)\equiv\theta_{i}$ (for a total of 4 real initial conditions, as required for a second-order complex equation of motion), the temporal evolution of $\rho_{0}(t)$, $\theta_{0}(t)$ must be obtained numerically; see Fig. 1 for some examples. As expected, because the effective potential is neither Coulombic nor of harmonic-oscillator form, the evolution of $\Phi_{0}$ does not form a closed curve in complex field space; in the absence of perturbations, it would eventually densely explore the entire complex field region that lies between the extremes of the $\rho_{0}$ motion.
@@ 2.2.1 Governing Equations
This is now a system of coupled oscillators with a time-dependent squared-frequency matrix $[\Omega(t)]^{2}$. Of course, full understanding of this system can only be obtained by evaluating $[\Omega(t)]^{2}$ on the numerical background solution and then solving the resulting coupled system numerically.
@@
where diagonalization by conjugation with the time-dependent $SO(2)$ rotation matrices [±$\mathcal{R}[\pm\theta_{0}(t)]$ gives rise to the matrix $\mathcal{D}$ whose diagonal entries are the time-dependent eigenvalues of $[\Omega(t)]^{2}$. From this, we note that (20) can be re-written as
@@
Because both the rotation matrices [±$\mathcal{R}[\pm\theta_{0}(t)]$ and the entries of $\mathcal{D}$ are time-dependent, we cannot however read off the full solution to (23) in any closed form from this information.
@@ 2.2.2 Simplified System Analytics at Small $\mathcal{J}_{0}$
$\Rightarrow\zeta_{0}(\tilde{t})$ $\approx\bar{\zeta}_{0}\cos(\sqrt{2}\cdot\tilde{t}+\varphi_{\zeta})\qquad\left[\zeta_{0}\ll 1,\>\mathcal{J}_{0}\ll\mu\rho_{0}^{2}\right]\>,$ (25)
@@
We have thus been able to cast the approximate EoM for the perturbations into the form $\partial_{t^{\prime}}^{2}f+[\omega^{\prime}_{i}]^{2}f=0$ where $[\omega_{i}^{\prime}]^{2}\equiv\omega_{0,i}^{2}+2\epsilon_{i}\cos(\sqrt{2}(t^{\prime}-t^{\prime}_{0}))$, otherwise known as the Mathieu equation, [^11] This is a slightly non-standard form of the Mathieu equation; it can easily be re-cast to the standard forms by means of rescaling the time variable and the parameters. which famously exhibits both narrow and broad parametric resonance instability phenomena, on which a vast body of literature exists (see, e.g, MagnusWinkler ; Kofman:1997yn ; Kovacic:2018weg ). We undertake a detailed analysis of the resulting Fourier-mode instabilities in App. C.
@@
As the field excursion $\bar{\zeta}_{0}$ increases from very small values, the $\tilde{k}$ ranges that define the narrow resonance bands where $\delta\rho$ and $\delta\sigma$ are unstable will grow in size and eventually merge when the upper limit of the $\delta\rho$ resonance band hits the lower limit of the $\delta\sigma$ band. This happens when $\sqrt{15/2}\cdot\bar{\zeta}_{0}=\sqrt{1/2-\bar{\zeta}_{0}}\Rightarrow\bar{\zeta}_{0}=1/5=0.2$, with the merger occurring at $\tilde{k}=\sqrt{3/10}\sim 0.55$. Note that this conclusion is marginally inconsistent with the $\delta\sigma$ resonance remaining narrow ($\bar{\zeta}_{0}\lesssim 1/6$, as discussed in App. C), but we expect that this analysis should still be approximately correct. For larger field excursions, we expect that ranges of $\tilde{k}$ will exist in which both $\delta\rho$ and $\delta\sigma$ will simultaneously parametrically resonant (possibly in a mix of narrow and broad resonances).
@@ 2.4 Fully Nonlinear Simulations
Finally, we complete the analysis of the parametric resonance by performing time-domain numerical simulations of the full nonlinear system (4). To that end, we discretize the classical field equation (4) in Cartesian coordinates on a two- or three-dimensional spatial grid covering a box of side length $2L$ and impose periodic boundary conditions. The initial conditions, $\Phi_{i}(\bm{x})$ and $\dot{\Phi}_{i}(\bm{x})$, are chosen to describe phase-coherent field configurations with average radial displacement $\rho_{i}$ and non-vanishing, spatially homogeneous charge density $\mathcal{J}_{0}$. This average field configuration is supplemented by a scale-invariant Gaussian random field $\mathcal{N}(\bm{x})$ with variance $\sigma_{\rho_{i}}$ and vanishing mean. Ultimately, this implies [^12] These initial conditions are of course slightly tuned in that we have assumed that all field values start with the same phase, but with different radial field values and initial angular speeds, such that $\mathcal{J}_{0}$ is homogeneous. More general perturbed initial conditions would also consider an initial small perturbation to the phase (i.e., the initial $\Phi_{i}$ field values would be sampled in a ball of complex field values centered on $\rho_{i}$). Because of nonlinearities and the different angular speeds initially imposed, we expect such phase differences to appear naturally from these initial conditions after a short amount of temporal evolution. Overall, this slightly tuned initial conditions is unlikely to yield qualitatively different results from the more general case. $\Phi_{i}(\bm{x})=\rho_{i}+\mathcal{N}(\bm{x})$ and $\dot{\Phi}_{i}(\bm{x})=i\mathcal{J}_{0}[\rho_{i}+\mathcal{N}(\bm{x})]^{-1}$. While our simulations are performed in the two real variables $\text{Re}[\Phi(t,\bm{x})]$ and $\text{Im}[\Phi(t,\bm{x})]$ (and hence are robust against issues associated with the ill-defined multi-valued angular degree of freedom at the location of vortices, where $|\Phi|\rightarrow 0$), we do also find it convenient to reconstruct the radial $\rho(t,\bm{x})\equiv(\text{Re}[\Phi(t,\bm{x})]^{2}+\text{Im}[\Phi(t,\bm{x})]^{2})^{1/2}$ and angular $\theta(t,\bm{x})\equiv\arctan(\text{Im}[\Phi(t,\bm{x})]/\text{Re}[\Phi(t,\bm{x})])\in[-\pi,\pi)$ degrees of freedom.
[…]
@@ 3.1 Global-String Profile
In two spatial dimensions, an isolated (anti-)vortex at $r=0$ in an otherwise unexcited system has a field profile $\Phi(\bm{x})=\rho(r)e^{\pm i\theta}$ with boundary conditions $\rho(0)=0$ and $\rho(\infty)=v$, where the $\pm$ sign choice specifies positive (vortex) or negative (anti-vortex) vorticity. Note that $\theta$ is multi-valued at $r=0$, and $\partial_{i}\theta$ diverges as a result. The radial profile of a vortex has been solved explicitly; see Weinberg:1992hc and references within. Schematically, the string has a ‘core’ region in which $\rho(r)$ increases from 00 to $\mathcal{O}(v)$ over a distance of $\mathcal{O}(1/\mu)$.
@@
The vortices in our analysis differ from isolated vortices in an otherwise-unexcited system in two major ways. First, for most of our simulations, our vortices reside in a background-field configuration for which the radial mode is excited (i.e., displaced significantly from the minimum of the potential); specifically, the average radial-mode displacement is $\rho>v$ in the counter-rotating regions and $\rho<v$ in the domain-boundary regions. Second, the background-field configuration in which the vortices reside can have $\dot{\theta}$ and/or $\partial_{i}{\theta}$ significantly different from zero. Both of these differences modify the vortex structure in ways that are analytically intractable to compute. Luckily, however, their topological nature still makes vortices or strings easy to identify, even absent knowledge of their spatial profile: they are simply points (in 2D) or lines (in 3D) where $\rho=0$ and around which $\theta$ changes by ±$\pm 2\pi$, even on an arbitrarily small loop. As a result, it is possible to count the number of vortices (and anti-vortices) in 2D, or measure the length of strings in 3D.

### v2

@@ 1 Introduction and Summary
The formation of topological defects in theories with global or gauged $U(1)$ symmetries has been extensively studied in both high-energy and condensed-matter contexts Kibble:1976sj ; Kibble:1980mv ; Zurek:1996sj . The Kibble–Zurek mechanism Kibble:1976sj ; Zurek:1993ek ; Zurek:1985qw is a well-known example in thermal systems, in which topological defects form as the early Universe cools below the critical temperature for spontaneous symmetry breaking. Such defects can have important phenomenological consequences (for instance for axion dark matter, where details are still under active investigation; see, e.g., Gorghetto:2020qws ; Buschmann:2021sdq ; Benabou:2023ghl ; Benabou:2023npn ; Gorghetto:2023vqu ; Benabou:2024msj ; Gorghetto:2024vnp for a selection of recent work). Formation of cosmic strings can also occur non-thermally as a result of quantum fluctuations during inflation Linde:1990yj ; PhysRevD.44.340 ; Vishniac:1986sk ; Lyth:1992tw . Defect formation during preheating scenarios has also been studied extensively Kasuya:1997ha ; Kasuya:1998td ; Kasuya:1999hy . After inflation, nonlinearities may drive the formation of topological defects in a large background field Abrikosov:1956sx ; osti_4551407 ; Galaiko1966FormationOV ; East:2022ppo ; East:2022rsi ; Kofman:1997yn ; Kawasaki:2013iha . In the gauged-$U(1)$ case (i.e., the Abelian Higgs model), in a system where the gauge field strength grows East:2022ppo ; East:2022rsi (either in the early Universe or via black-hole superradiance) and surpasses a critical threshold, a superheated phase transition occurs and a dense network of strings is produced. [^1] We will use the terms ‘vortices’ and ‘strings’ when discussing the dynamics presented in this paper. Although vortices are generally associated with $(2+1)$-dimensional systems and strings with $(3+1)$-dimensional systems, we use them interchangeably, depending on the context. For all practical purposes of this paper, there is no difference between these two terms. This string network absorbs energy from the background gauge fields, grows in size, and persists in the system. These dynamics resemble the vortex-forming phase transition in a superconductor, which is described in the framework of the Ginzburg–Landau model Ginzburg:1950sr ; Abrikosov:1956sx (see App. A for a short review of the relevant literature).
@@
The paper is organized as follows. In Sec. 2, we begin with a purely analytical perturbative analysis valid for small $U(1)$ charge density that demonstrates the existence of a parametric resonance instability Tkachev:1998dc for spatial field perturbations $\delta\Phi$ around a spatially homogeneous background field evolution $\Phi_{0}$ [Secs. 2.1 and 2.2]. We extend this perturbative analysis to larger $U(1)$ charge densities via a numerical study of the linearized system of equations governing the evolution of the perturbations $\delta\Phi$ Co:2020dya , demonstrating that this parametric resonance remains present in this case [Sec. 2.3]. We then analyze the fully non-linear evolution of the $\Phi$ field from the epoch of the initial parametric resonance through to the onset of the formation of the counter-rotating regions (CRR) that emerge when the perturbations grow beyond the linear regime phase-coherently over a range of momentum modes [Sec. 2.4]; we do this both for small and, for the first time, large $U(1)$ charge densities. Finally, we summarize the main points of this section [Sec. 2.5].
@@ 2 Perturbation Growth to the Onset of Vortex Formation
In this section we show that a spatially homogeneous scalar field subject to a wine-bottle potential, initially either with or without field-space angular momentum, is not a stable field configuration: rather, it exhibits (classical) parametric instabilities that lead to growth of spatial inhomogeneities. Some of the results presented in this section were previously discussed in, e.g., Dolgov:1989us ; Traschen:1990sw ; Kofman:1994rk ; Kofman:1997yn ; Co:2020dya .
For now, we ignore Hubble expansion and work in $(D+1)$-dimensional Minkowski spacetime (mostly minus convention; signature 1D1-D). The Lagrangian (1) yields the equations of motion (EoM)
@@ 2.1 Background Field Evolution
Although the background-field trajectory $\Phi_{0}(t)$ is completely specified by a choice of the two conserved quantities $\mathcal{J}_{0}$ and $\mathcal{E}_{0}$, supplemented with an initial field point on the trajectory fixed by $\rho_{0}(0)\equiv\rho_{i}$ and $\theta_{0}(0)\equiv\theta_{i}$ (for a total of 4 real initial conditions, as required for a second-order complex equation of motion), the temporal evolution of $\rho_{0}(t)$, $\theta_{0}(t)$ must be obtained numerically (but see, e.g., Appendix G of Gouttenoire:2021jhk for semi-analytical solutions valid in certain limits); see Fig. 1 for some examples. As expected, because the effective potential is neither Coulombic nor of harmonic-oscillator form, the evolution of $\Phi_{0}$ does not form a closed curve in complex field space; in the absence of perturbations, it would eventually densely explore the entire complex field region that lies between the extremes of the $\rho_{0}$ motion.
@@ 2.2.1 Governing Equations
This is now a system of coupled oscillators with a time-dependent squared-frequency matrix $\Omega(t)]^{2}$. Of course, full understanding of this system can only be obtained by evaluating $\Omega(t)]^{2}$ on the numerical background solution and then solving the resulting coupled system numerically.
@@
where diagonalization by conjugation with the time-dependent $SO(2)$ rotation matrices $\mathcal{R}[\pm\theta_{0}(t)]$ gives rise to the matrix $\mathcal{D}$ whose diagonal entries are the time-dependent eigenvalues of $\Omega(t)]^{2}$. From this, we note that (20) can be re-written as
@@
Because both the rotation matrices $\mathcal{R}[\pm\theta_{0}(t)]$ and the entries of $\mathcal{D}$ are time-dependent, we cannot however read off the full solution to (23) in any closed form from this information.
@@ 2.2.2 Simplified System Analytics at Small $\mathcal{J}_{0}$
[…]
@@
We have thus been able to cast the approximate EoM for the perturbations into the form $\partial_{t^{\prime}}^{2}f+[\omega^{\prime}_{i}]^{2}f=0$ where $\omega_{i}^{\prime}]^{2}\equiv\omega_{0,i}^{2}+2\epsilon_{i}\cos(\sqrt{2}(t^{\prime}-t^{\prime}_{0}))$, otherwise known as the Mathieu equation, [^11] This is a slightly non-standard form of the Mathieu equation; it can easily be re-cast to the standard forms by means of rescaling the time variable and the parameters. which famously exhibits both narrow and broad parametric resonance instability phenomena, on which a vast body of literature exists (see, e.g., Dolgov:1989us ; Traschen:1990sw ; Kofman:1994rk ; Kofman:1997yn ; MagnusWinkler ; Kovacic:2018weg ). We undertake a detailed analysis of the resulting Fourier-mode instabilities in App. C.
@@
As the field excursion $\bar{\zeta}_{0}$ increases from very small values, the $\tilde{k}$ ranges that define the narrow resonance bands where $\delta\rho$ and $\delta\sigma$ are unstable will grow in size and eventually merge when the upper limit of the $\delta\rho$ resonance band hits the lower limit of the $\delta\sigma$ band. This happens when $\sqrt{15/2}\cdot\bar{\zeta}_{0}=\sqrt{1/2-\bar{\zeta}_{0}}\Rightarrow\bar{\zeta}_{0}=1/5=0.2$, with the merger occurring at $\tilde{k}=\sqrt{3/10}\sim 0.55$. Note that this conclusion is marginally inconsistent with the $\delta\sigma$ resonance remaining narrow ($\bar{\zeta}_{0}\lesssim 1/6$, as discussed in App. C), but we expect that this analysis should still be approximately correct. For larger field excursions, we expect that ranges of $\tilde{k}$ will exist in which both $\delta\rho$ and $\delta\sigma$ will be simultaneously parametrically resonant (possibly in a mix of narrow and broad resonances).
@@ 2.4 Fully Nonlinear Simulations
Finally, we complete the analysis of the parametric resonance by performing time-domain numerical simulations of the full nonlinear system (4). To that end, we discretize the classical field equation (4) in Cartesian coordinates on a two- or three-dimensional spatial grid covering a box of side length $2L$ and impose periodic boundary conditions. The initial conditions, $\Phi_{i}(\bm{x})$ and $\dot{\Phi}_{i}(\bm{x})$, are chosen to describe phase-coherent field configurations with average radial displacement $\rho_{i}$ and non-vanishing, spatially homogeneous charge density $\mathcal{J}_{0}$. This average field configuration is supplemented by a scale-invariant Gaussian random field $\mathcal{N}(\bm{x})$ with variance $\sigma_{\rho_{i}}$ and vanishing mean. Ultimately, this implies [^12] These initial conditions are of course slightly tuned in that we have assumed that all field values start with the same phase, but with different radial field values and initial angular speeds, such that $\mathcal{J}_{0}$ is homogeneous. More general perturbed initial conditions would also consider an initial small perturbation to the phase (i.e., the initial $\Phi_{i}$ field values would be sampled in a ball of complex field values centered on $\rho_{i}$). Because of nonlinearities and the different angular speeds initially imposed, we expect such phase differences to appear naturally from these initial conditions after a short amount of temporal evolution. Overall, these slightly tuned initial conditions are unlikely to yield qualitatively different results from the more general case. $\Phi_{i}(\bm{x})=\rho_{i}+\mathcal{N}(\bm{x})$ and $\dot{\Phi}_{i}(\bm{x})=i\mathcal{J}_{0}[\rho_{i}+\mathcal{N}(\bm{x})]^{-1}$. While our simulations are performed in the two real variables $\text{Re}[\Phi(t,\bm{x})]$ and $\text{Im}[\Phi(t,\bm{x})]$ (and hence are robust against issues associated with the ill-defined multi-valued angular degree of freedom at the location of vortices, where $|\Phi|\rightarrow 0$), we do also find it convenient to reconstruct the radial $\rho(t,\bm{x})\equiv(\text{Re}[\Phi(t,\bm{x})]^{2}+\text{Im}[\Phi(t,\bm{x})]^{2})^{1/2}$ and angular $\theta(t,\bm{x})\equiv\arctan(\text{Im}[\Phi(t,\bm{x})]/\text{Re}[\Phi(t,\bm{x})])\in[-\pi,\pi)$ degrees of freedom.
The addition of the Gaussian [^13] The initial conditions we assume are of course scale-invariant, but this is likely not strictly necessary. Recall that the parametric resonance instabilities drive exponential growth of the unstable modes. So long as the unstable modes are not exponentially suppressed in the initial conditions, they will thus still grow to dominate the perturbation spectrum in short order, and we would expect similar results, even if the initial perturbation spectrum is not white noise (so long as all initial perturbation modes are still small). Note however that while our specific initial conditions are inspired by similar initial conditions observed in the axion kinetic misalignment / kination literature, we do not specify a microphysical model for them here. random field $\mathcal{N}(\bm{x})$ ensures that all unstable spatial Fourier modes are excited with (roughly) the same initial amplitude and the same initial complex phase. [^14] On a numerical grid, the finite box size and spatial resolution impose lower and upper bounds, respectively, on resolved Fourier modes. We ensure that the fastest growing (i.e., most relevant) modes are well-resolved in all cases considered. See App. B for further details. The evolution of the system from the initial conditions follows the background evolution, discussed in Sec. 2.1, for sufficiently small $\sigma_{\rho_{i}}$ at early times. That is, early on in the evolution, $\langle\rho(t,\bm{x})\rangle\approx\rho_{0}(t)$ and $\langle\theta(t,\bm{x})\rangle\approx\theta_{0}(t)$ as defined in the previous section. Here, we denote the spatial average by $\langle\ldots\rangle$; furthermore, for brevity, we drop the explicit spatial and time dependencies of $\rho(t,\bm{x})\rightarrow\rho$ and $\theta(t,\bm{x})\rightarrow\theta$ unless stated otherwise. The energy densities of these respective degrees of freedom,
@@ 3.1 Global-String Profile
In two spatial dimensions, an isolated (anti-)vortex at $r=0$ in an otherwise unexcited system has a field profile $\Phi(\bm{x})=\rho(r)e^{\pm i\theta}$ with boundary conditions $\rho(0)=0$ and $\rho(\infty)=v$, where the $\pm$ sign choice specifies positive (vortex) or negative (anti-vortex) vorticity. Note that $\theta$ is multi-valued at $r=0$, and $\partial_{i}\theta$ diverges as a result. The radial profile of a vortex has been solved explicitly; see Weinberg:1992hc and references within. Schematically, the string has a ‘core’ region in which $\rho(r)$ increases from 0 to $\mathcal{O}(v)$ over a distance of $\mathcal{O}(1/\mu)$.
@@
The vortices in our analysis differ from isolated vortices in an otherwise-unexcited system in two major ways. First, for most of our simulations, our vortices reside in a background-field configuration for which the radial mode is excited (i.e., displaced significantly from the minimum of the potential); specifically, the average radial-mode displacement is $\rho>v$ in the counter-rotating regions and $\rho<v$ in the domain-boundary regions. Second, the background-field configuration in which the vortices reside can have $\dot{\theta}$ and/or $\partial_{i}{\theta}$ significantly different from zero. Both of these differences modify the vortex structure in ways that are analytically intractable to compute. Luckily, however, their topological nature still makes vortices or strings easy to identify, even absent knowledge of their spatial profile: they are simply points (in 2D) or lines (in 3D) where $\rho=0$ and around which $\theta$ changes by $\pm 2\pi$, even on an arbitrarily small loop. As a result, it is possible to count the number of vortices (and anti-vortices) in 2D, or measure the length of strings in 3D.